OpenStudy (lgbasallote):

OpenStudy user a64 has twice the fans OpenStudy user L has. If L gains 110 fans per month and a64 gains 50 fans per month, how long will it take for L to get twice the fans a64 has? is this solveable now?

OpenStudy (anonymous):

Apply what I said before.

OpenStudy (anonymous):

Let a64 fans be x and L fans be y. So, x = 2y y + 110 = 2(x + 50).. Am I doing right or not??

OpenStudy (lgbasallote):

hmm

OpenStudy (anonymous):

@waterineyes, I answered all of this on lgba's previous question. I just don't think he read it all.

OpenStudy (lgbasallote):

i read it..i just need to practice it

OpenStudy (lgbasallote):

reading wont suffice after all will it

OpenStudy (anonymous):

Tell me you both of you, I am right in doing that or not??

OpenStudy (lgbasallote):

so y + 110 = 2(x + 50) y + 110 = 2(2y + 50) y + 110 = 4y + 100 10 = 3y y = 3.33 months so it's saying that in 3 months i'll have double the fans of amistre?

OpenStudy (lgbasallote):

did i do the solution right @Limitless ?

OpenStudy (anonymous):

@lgbasallote, just a moment.

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

\(L=L_i+110m\) \(A=2L_i+50m\) \(L=2A\) \(L_i+110m=4L_i+100m\) \(10m=3L_i\) \(m=\frac{3L_i}{10}\) \(L_i\) is the initial fans you have.

OpenStudy (lgbasallote):

oh woops i forgot to put variables hehe

OpenStudy (lgbasallote):

3/10 is like 0.3 right?

OpenStudy (lgbasallote):

that doesnt sound right... o.O